OpenStudy (anonymous):

how to combine like terms 8r-(3s-5r) how do i get the answer

OpenStudy (asnaseer):

first deal with the brackets

OpenStudy (asnaseer):

-(3s-5r) = ?

OpenStudy (anonymous):

i thought you couldnt solve it like the brackets because it has a s and a r ? and i dont know how to solve that

OpenStudy (asnaseer):

you need to distribute the negative sign outside the brackets into each term within the brackets

OpenStudy (asnaseer):

are you aware of these rules: 1. minus and minus makes a plus 2. minus and plus makes a minus 3. plus and minus makes a minus 4. plus and plus makes a plus

OpenStudy (asnaseer):

e.g. -(3) = -3 -(-3) = 3

OpenStudy (anonymous):

yeah i undersatnd thtt but im not understandng how to solve it

OpenStudy (asnaseer):

so use these rules to remove the brackets around this: -(3s-5r) = ?

OpenStudy (anonymous):

thxs

OpenStudy (asnaseer):

do you understand?

OpenStudy (anonymous):

yeaa

OpenStudy (asnaseer):

ok - so what do you get after removing the brackets?

OpenStudy (asnaseer):

this may /seem/ like a painful process but if you stick to it for this problem then you will find all other similar problems easy to solve.

OpenStudy (asnaseer):

I can explain in more detail if you are willing to learn?